Inessential Refrigerator Restocking

For Morons

(Revision: 1.1 )

Dave Cho (dcctdw@athena.mit.edu)

May 9, 2005

Contents

i

1 Introduction: What is a refrigerator?

From Inessential Refrigerator Restocking :

The SIPB owns a refrigerator for the express purpose of cooling sodas and juices. It is usually appreciated by people that their sodas and juice drinks are cold.

Now, the big white rectangular object in the rear of the SIPB office, across from the stereo, is the refrigerator. In this very same refrigerator are kept sodas and juices. There is a very high probability that you even get a soda or juice from this refrigerator at semi-regular intervals.

2 Getting Started

From Inessential Refrigerator Restocking :

When you see that the refrigerator is low on a particular item (for instance, Coca-Cola, also known as Coke), then you should pause and think. We realize this is difficult for some people.

Silly us, we thought that thinking was a procedure that most people could grasp. We shall go into this at more length.

To think means to ponder, to consider, to reason. It is widely believed that MIT students are very good at thinking. This means that if you aren’t good at thinking, you should consider autodefenestration.

3 What to Start With

Look in the refrigerator. This means that when you open the door, look in. Note that this momentous event probably happens as you reach for a soda or a juice. Now, savor this momentous happening by looking further!. Do you notice that there are two or three of that soda or juice left? Eureka! Oh, frabjous day, we have a miraculous breakthrough!

4 Processing Your Discovery

From Inessential Refrigerator Restocking :

At this point, you should inspect the stack of sodas and juices located next to the music equipment.

In Inessential Refrigerator Restocking , we didn’t explain this well enough. What this means is that you should push onto your operations stack your current mindset. (The author realizes that many platforms cannot support this. The author suggests that those platforms have linear acceleration experiments performed on them, as that is the only possible way to get any sort of useful societal contribution from aforementioned platforms.) Now, get another case of the appropriate soda or juice and put it inside the refrigerator. See Inessential Refrigerator Restocking for more details.

5 Wait! Did you understand that? Probably not, so let’s try again...

The author shall now type slower for those who didn’t understand.

• Take a longer look while your head is cooling off inside the refrigerator

• Determine if you’re taking one of the last few cans.

• Take a few minutes to do something useful about it.

If you still don’t understand this, please follow the instructions contained in the above parenthetical remark about useless platforms.

6 One last note

There are probably many SIPB members who can help the novice with this new and terrifying concept. Feel free to ask for assistance.

7 Where to Get More Information

From Inessential Refrigerator Restocking :

Try calling your mother, and asking her about common courtesy. She’ll probably be happy to explain.

If this doesn’t work, try stapling bagels to your face.

8 Acknowledgments

The author would like to thank Amazingly Stoopidicus Maximus and Stoopid Cluless Dumfuck, whose existence necessitated the creation of this document.